# Iterating $\mathsf{PP}$-packages without Choice: A Cohen symmetric seed and a localization framework

### v3

@@ 4.2 The $\mathsf{AC}_{\mathsf{WO}}$-package forcing $\mathbb{R}_{f}$
Fix a transitive model $M\models\mathsf{ZF}+\mathsf{DC}$ extending the Cohen seed $\mathcal{N}$ (again, in practice $M$ will be an intermediate stage of our iteration). Retain the fixed seed parameter $S$ from Remark 4.4.
@@
Assuming $\mathsf{SVC}(S)$, Ryan–Smith shows that $\mathsf{AC}_{\mathsf{WO}}$ is equivalent to the bounded scheme $\mathsf{AC}_{<\aleph^{*}(S)}$ (more precisely, $\mathsf{AC}_{<\aleph^{*}(S)}(S)$ in his notation); see [17, Cor. 3.3]. For fixed $\lambda$, the corresponding “splitting” reformulations of $\mathsf{AC}_{\lambda}$ appear in [17, Prop. 3.2]. Thus, in the presence of $\mathsf{SVC}(S)$ it suffices to schedule splittings only for surjections onto ordinals $\lambda<\aleph^{*}(S)$.
@@
Unlike the $\mathsf{PP}$-packages $\mathbb{Q}_{f}$, which are explicitly localized to $X,Y\subseteq T=\mathcal{P}(S)$, the $\mathsf{AC}_{\mathsf{WO}}$-package must be global in nature: $\mathsf{AC}_{\mathsf{WO}}$ speaks about all well-ordered index sets. The point of Remark 4.22 is that, assuming $\mathsf{SVC}(S)$, it is enough to schedule splittings for surjections onto $\lambda<\aleph^{*}(S)$; this reduction is precisely what makes the bookkeeping manageable.
@@
The $\omega_{1}$-closure of $\mathbb{R}_{f}$ is proved inside $M$ using $\mathsf{DC}$ (equivalently $AC_{\omega}$), via “countable union of countable sets is countable”. In bare $\mathsf{ZF}$ this need not hold.
@@
Apply Lemma 4.18 with $\mathbb{P}=\mathbb{R}_{f}$. Countable closure holds by Lemma 4.27. ∎
@@ 5 Iterating the packages over $\mathcal{N}$
This section describes a class-length countable-support symmetric iteration built to force local splitting principles while preserving $\mathsf{DC}$ and killing $\mathsf{AC}$. Here is the dependency flow.
@@
We now describe the symmetric iteration that forces the two conjuncts $\mathsf{PP}^{\mathrm{split}}\!\restriction T$ and $\mathsf{AC}_{\mathsf{WO}}$ over the Cohen symmetric seed model $\mathcal{N}$, while preserving $\neg\mathsf{AC}$ (witnessed by the non-well-orderability of $A$). All technical iteration infrastructure is imported via the API of Remark 2.1. The purpose of this section is to isolate the package systems, the bookkeeping targets, and the verification that the final symmetric model satisfies the desired principles.
@@
The argument is deliberately modular: the seed model technology (Section 3), the localized $\mathsf{PP}$-packages (Section 4.1), the $\mathsf{AC}_{\mathsf{WO}}$-packages (Section 4.2), and the symmetric-iteration “API” (Remark 2.1) are separated so that if a hypothesis later needs to be strengthened or a package definition adjusted, it can typically be done locally without rewriting unrelated parts of the paper.
@@ 5.1 Package symmetric systems at a stage
Fix a transitive $M\models\mathsf{ZF}+\mathsf{DC}$ extending $\mathcal{N}$ (in practice $M=M_{\alpha}$ at an intermediate stage of the iteration). Keep the parameters $S$ and $T=\mathcal{P}(S)$ as in Remark 4.4.
Remark 5.3 (An ambient $\operatorname{Sym}(\omega_{1})$-action on $T=\mathcal{P}(S)$).
We will often use the ambient action of $\operatorname{Sym}(\omega_{1})$ on the stage-0 seed parameter $S=A^{\omega}$ and on $T=\mathcal{P}(S)$ by permuting the Cohen reals $c_{\alpha}\in A$. Concretely, for $\pi\in\operatorname{Sym}(\omega_{1})$ we set $\pi\cdot c_{\alpha}:=c_{\pi(\alpha)}$, extend pointwise to sequences $s\in A^{\omega}$, and then define $\pi\cdot X:=\{\pi\cdot s:s\in X\}$ for $X\subseteq S$, i.e. $\pi\cdot X\in T$. We regard this as an action which fixes all ordinals pointwise; i.e. $\pi\cdot\xi=\xi$ for every ordinal $\xi$.
We furthermore extend this action to every set $x$ with $\operatorname{tc}(x)\subseteq V(T)$ by recursion on rank:
$\pi\cdot x\;:=\;\{\pi\cdot y:y\in x\}.$
In particular, if $f:X\to Y$ is a function with $\operatorname{tc}(f)\subseteq V(T)$ (e.g. $X,Y\subseteq T$), we identify $f$ with its graph and set
@@
Let $\alpha$ be any stage and let $k\in\ker(\pi^{0}_{\alpha})$. Then $k$ fixes $A$, $S=A^{\omega}$, and $T=\mathcal{P}(S)$ pointwise for the ambient $\operatorname{Sym}(\omega_{1})$-action of Remark 5.3. Consequently, $k$ fixes every set/function whose transitive closure is contained in $V(T)\cup\mathrm{Ord}$; in particular, if $f:X\to Y$ with $X,Y\subseteq T$, or $f:S\times\eta\twoheadrightarrow\lambda$ with $\eta,\lambda\in\mathrm{Ord}$, then $k\cdot f=f$.
@@
Put $\sigma:=\pi^{0}_{\alpha}(k)\in\operatorname{Sym}(\omega_{1})$. Since $k\in\ker(\pi^{0}_{\alpha})$ we have $\sigma=\operatorname{id}$. By Remark 5.3, the ambient action on $A$, on $S=A^{\omega}$, and on $T=\mathcal{P}(S)$ is induced from the action $\alpha\mapsto\sigma(\alpha)$ on indices of the Cohen reals. Therefore $\sigma=\operatorname{id}$ implies that $k$ fixes $A$, and hence also $S$ and $T$, pointwise.
@@
Lemma 5.6 (Orbit packages are $\omega_{1}$-closed and preserve $\mathsf{DC}$).
Assume $M\models\mathsf{ZF}+\mathsf{DC}$ and let $\mathbb{S}_{[f]}$ be either $\mathbb{Q}_{[f]}$ or $\mathbb{R}_{[f]}$. Then:
  1. (i)
    $\mathbb{S}_{[f]}$ is $\omega_{1}$-closed in $M$;
  2. (ii)
    $\mathbb{S}_{[f]}$ adds no new $\omega$-sequences of ordinals (in particular, no new reals);
  3. (iii)
    forcing with $\mathbb{S}_{[f]}$ over $M$ preserves $\mathsf{DC}$.
@@
(i) Let $\langle\vec{p}_{n}:n<\omega\rangle$ be decreasing in $\mathbb{S}_{[f]}$. For each $g\in[f]$ define $q_{g}:=\bigcup_{n<\omega}p_{n,g}$, where $p_{n,g}$ is the $g$-coordinate of $\vec{p}_{n}$ (taking $p_{n,g}=\emptyset$ if $g\notin\operatorname{supp}(\vec{p}_{n})$). Since each factor $\mathbb{Q}_{g}$ is $\omega_{1}$-closed by Lemma 4.15 (and each factor $\mathbb{R}_{g}$ by Lemma 4.27), $q_{g}$ is a condition in that factor. Let $\vec{q}$ have support $\operatorname{supp}(\vec{q}):=\bigcup_{n<\omega}\operatorname{supp}(\vec{p}_{n})$ and coordinates $q_{g}$ for $g\in\operatorname{supp}(\vec{q})$. Because $M\models\mathsf{DC}$ (hence $\mathsf{AC}_{\omega}$), a countable union of countable sets is countable, so $\operatorname{supp}(\vec{q})$ is countable. Thus $\vec{q}\in\mathbb{S}_{[f]}$ and $\vec{q}\leq\vec{p}_{n}$ for all $n$.
(ii) and (iii) now follow by applying Lemma 4.18 to $\mathbb{P}=\mathbb{S}_{[f]}$, using (i) for countable closure. ∎
@@
At limit stages, $\omega_{1}$-completeness (and normality) of the limit filter is provided by the iteration machinery (API item (d), cf. [5, Thm. 5.5]).
@@
Lemma 5.9 (Countable support of package coordinates).
For every package coordinate $d$ at any successor stage $\alpha+1$ of the iteration, there exists a countable set $E_{d}\subseteq\omega_{1}$ such that $\operatorname{Fix}(E_{d})\leq\operatorname{Stab}(d)$, where $\operatorname{Stab}(d)=\{\pi\in\operatorname{Sym}(\omega_{1}):\pi\cdot d=d\}$ is the stabilizer of $d$ as an element of its orbit package.
@@
We proceed by induction on stages. At stage 0 there are no package coordinates, so the statement is vacuous.
At successor stage $\alpha+1$, package coordinates are elements of orbit packages $f]=\{\pi\cdot f:\pi\in\operatorname{Sym}(\omega_{1})\}$, where $f$ is a surjection in the symmetric model $M_{\alpha}$.
Step 1: The seed surjection $f$ has countable support.
Since $f\in M_{\alpha}$, there is a hereditarily symmetric $\mathbb{P}_{\alpha}$-name $\dot{f}$ with $f=\dot{f}^{G_{\alpha}}$. By the $\Delta$-support equivalence at stage $\alpha$ (which holds by the inductive hypothesis that the Task 1 filter infrastructure is established at stage $\alpha$), there exists a countable admissible $(E_{f},D_{f})$ such that
$\Delta_{\alpha}^{\uparrow}(E_{f},D_{f})\leq\operatorname{Stab}_{\mathscr{G}_{\alpha}}(\dot{f}).$
Step 2: Bridge from $\mathscr{G}_{\alpha}$-stabilizer to $\operatorname{Sym}(\omega_{1})$-stabilizer.
We claim $\operatorname{Fix}(E_{f})\leq\operatorname{Stab}(f)$. Let $\pi\in\operatorname{Fix}(E_{f})$.
By Lemma 5.43 (applied at stage $\alpha$, valid by the inductive hypothesis), there exists a diagonal lift $\widehat{\pi}^{\,D_{f}}\in\mathscr{G}_{\alpha}$ with $\pi^{0}_{\alpha}(\widehat{\pi}^{\,D_{f}})=\pi$ and which fixes every coordinate in $D_{f}$ pointwise. By Definition 5.27, this implies $\widehat{\pi}^{\,D_{f}}\in\Delta_{\alpha}^{\uparrow}(E_{f},D_{f})\leq\operatorname{Stab}_{\mathscr{G}_{\alpha}}(\dot{f})$. Hence
$\widehat{\pi}^{\,D_{f}}\cdot\dot{f}=\dot{f}\quad\text{and therefore}\quad\widehat{\pi}^{\,D_{f}}\cdot f=f.$
Let $\widehat{\pi}^{\,\varnothing}\in\mathscr{G}_{\alpha}$ be the standard lift of $\pi$ (the diagonal lift with empty protection set), so $\pi^{0}_{\alpha}(\widehat{\pi}^{\,\varnothing})=\pi$. Set
$k:=(\widehat{\pi}^{\,\varnothing})^{-1}\cdot\widehat{\pi}^{\,D_{f}}\in\mathscr{G}_{\alpha}.$
Then $\pi^{0}_{\alpha}(k)=\operatorname{id}$, i.e. $k\in\ker(\pi^{0}_{\alpha})$. By Lemma 5.4, since $\operatorname{tc}(f)\subseteq V(T)\cup\mathrm{Ord}$, we have $k\cdot f=f$. Therefore,
$f=\widehat{\pi}^{\,D_{f}}\cdot f=\widehat{\pi}^{\,\varnothing}\cdot(k\cdot f)=\widehat{\pi}^{\,\varnothing}\cdot f.$
Finally, note that if $g,h\in\mathscr{G}_{\alpha}$ satisfy $\pi^{0}_{\alpha}(g)=\pi^{0}_{\alpha}(h)$, then $g^{-1}h\in\ker(\pi^{0}_{\alpha})$, so $g\cdot f=h\cdot f$ by Lemma 5.4. In particular, the action of the standard lift $\widehat{\pi}^{\,\varnothing}$ on $f$ depends only on $\pi=\pi^{0}_{\alpha}(\widehat{\pi}^{\,\varnothing})$, hence $\pi\cdot f=\widehat{\pi}^{\,\varnothing}\cdot f=f$. Hence
$\pi\cdot f=\widehat{\pi}^{\,\varnothing}\cdot f=f,$
and thus $\pi\in\operatorname{Stab}(f)$. This proves $\operatorname{Fix}(E_{f})\leq\operatorname{Stab}(f)$.
Step 3: Orbit elements inherit countable support.
For any $g=\sigma\cdot f\in[f]$ (where $\sigma\in\operatorname{Sym}(\omega_{1})$), we have
$\operatorname{Stab}(g)=\operatorname{Stab}(\sigma\cdot f)=\sigma\cdot\operatorname{Stab}(f)\cdot\sigma^{-1}.$
Since $\operatorname{Fix}(E_{f})\leq\operatorname{Stab}(f)$, conjugating gives
$\operatorname{Fix}(\sigma[E_{f}])=\sigma\cdot\operatorname{Fix}(E_{f})\cdot\sigma^{-1}\leq\sigma\cdot\operatorname{Stab}(f)\cdot\sigma^{-1}=\operatorname{Stab}(g).$
Thus $E_{g}:=\sigma[E_{f}]$ is a countable support for $g$. ∎
@@
The proof of Lemma 5.9 invokes Lemma 5.43 (existence of diagonal lifts) at stage $\alpha$. This does not create a circular dependency: the proof of Lemma 5.43 constructs diagonal lifts by direct recursion on stages using the package-level diagonal machinery of Section 5.1, and does not use Lemma 5.9, Lemma 5.40, or Lemma 5.50.
@@
Let $\mathscr{G}_{[f]}:=\operatorname{Aut}(\mathbb{S}_{[f]})$.

### v4

@@ 4.2 The $\mathsf{AC}_{\mathsf{WO}}$-package forcing $\mathbb{R}_{f}$
Fix a transitive model $M\models\mathsf{ZF}+\mathsf{DC}$ extending the Cohen seed $\mathcal{N}$ (again, in practice $M$ will be an intermediate stage of our iteration). Retain the fixed seed parameters $S,T$ from Definition 3.27 (and recall from Remark 4.4 that $T$ is kept fixed throughout the iteration).
@@
Assuming $\mathsf{SVC}(S)$, Ryan–Smith shows that $\mathsf{AC}_{\mathsf{WO}}$ is equivalent to the bounded scheme $\mathsf{AC}_{<\aleph^{*}(S)}$ (more precisely, $\mathsf{AC}_{<\aleph^{*}(S)}(S)$ in his notation); see [18, Cor. 3.3]. For fixed $\lambda$, the corresponding “splitting” reformulations of $\mathsf{AC}_{\lambda}$ appear in [18, Prop. 3.2]. Thus, in the presence of $\mathsf{SVC}(S)$ it suffices to schedule splittings only for surjections onto ordinals $\lambda<\aleph^{*}(S)^{M}$. In our application $\mathsf{SVC}(S)$ holds in each stage model; see Lemma 5.69 below.
@@
Unlike the localized $\mathsf{PP}$-packages $\mathbb{Q}_{f}$ (whose definitions quantify only over $X,Y\subseteq T$, with $T$ the fixed seed parameter from $\mathcal{N}$), the $\mathsf{AC}_{\mathsf{WO}}$-package is necessarily global: $\mathsf{AC}_{\mathsf{WO}}$ concerns all well-ordered index sets. The reduction in Remark 4.20 shows that, assuming $\mathsf{SVC}(S)$, it suffices to schedule splittings only for surjections onto ordinals $\lambda<\aleph^{*}(S)^{M}$, which is what makes the bookkeeping manageable.
@@
The countable closure of $\mathbb{R}_{f}$ (Lemma 4.25) is proved inside $M$ using $\mathsf{DC}$ (equivalently $\mathsf{AC}_{\omega}$), via “countable union of countable sets is countable”. In bare $\mathsf{ZF}$ this need not hold.
@@
Apply Lemma 4.16 with $\mathbb{U}=\mathbb{R}_{f}$. Countable closure holds by Lemma 4.25. ∎
@@ 5 Iterating the packages over $\mathcal{N}$
This section describes a class-length countable-support symmetric iteration built to force local splitting principles while preserving $\mathsf{DC}$ and maintaining $\neg\mathsf{AC}$.
Here is the dependency flow.
@@
We now describe the symmetric iteration that forces the two conjuncts $\mathsf{PP}^{\mathrm{split}}\!\restriction T$ and $\mathsf{AC}_{\mathsf{WO}}$ over the Cohen symmetric seed model $\mathcal{N}$, while maintaining $\neg\mathsf{AC}$ (witnessed by the non-well-orderability of $A$). All technical iteration infrastructure is imported via the API of Remark 2.4 (in particular, existence of the class-length iteration is used exactly under the hypotheses of [5, Theorem 3.25]). The purpose of this section is to isolate the package systems, the bookkeeping targets, and the verification that the final symmetric model satisfies the desired principles.
@@
The argument is deliberately modular: the seed model technology (Section 3), the localized $\mathsf{PP}$-packages (Subsection 4.1), the $\mathsf{AC}_{\mathsf{WO}}$-packages (Subsection 4.2), and the symmetric-iteration “API” (Remark 2.4) are separated so that if a hypothesis later needs to be strengthened or a package definition adjusted, it can typically be done locally without rewriting unrelated parts of the paper.
@@ 5.1 Package symmetric systems at a stage
Fix a transitive $M\models\mathsf{ZF}+\mathsf{DC}$ extending $\mathcal{N}$ (in practice $M=M_{\alpha}$ at an intermediate stage of the iteration). Keep the seed parameters $S$ and $T$ from Definition 3.27. Thus $S=A^{\omega}$, and $T=\mathcal{P}(S)^{\mathcal{N}}$ is computed in the seed model and then treated as a fixed parameter in all later stages (we do not re-interpret $T$ as $\mathcal{P}(S)^{M}$).
Remark 5.3 (An ambient $\mathscr{G}$-action on $S$ and $T$).
Recall $\mathscr{G}=\operatorname{Sym}(\omega_{1})$ (Definition 3.3). Using the Cohen reals $c_{\alpha}\in A$, we obtain an induced action of $\mathscr{G}$ on $A$ by $\pi\cdot c_{\alpha}:=c_{\pi(\alpha)}$, hence on $S=A^{\omega}$ by coordinatewise action, and on $\mathcal{P}(S)$ by $\pi\cdot X:=\{\pi\cdot s:s\in X\}$. Since $T=\mathcal{P}(S)^{\mathcal{N}}\subseteq\mathcal{P}(S)$, we use the restricted action on $T$.
We regard this as an action fixing all ordinals pointwise. More generally, whenever $x$ is a set with $\operatorname{tc}(x)\subseteq V(T)\cup\mathrm{Ord}$, we extend the action recursively by $\pi\cdot x:=\{\pi\cdot y:y\in x\}$. In particular, for a function $f:X\to Y$ with $\operatorname{tc}(f)\subseteq V(T)\cup\mathrm{Ord}$ (e.g. $X,Y\subseteq T\times\eta$ for an ordinal $\eta$), we identify $f$ with its graph and set
@@
Let $\alpha$ be any stage and let $k\in\ker(\pi^{0}_{\alpha})$, where $\pi^{0}_{\alpha}$ is the stage-0 projection of Definition 5.23. Then $k$ fixes $A$, $S=A^{\omega}$, and the fixed seed parameter $T=\mathcal{P}(S)^{\mathcal{N}}$ pointwise for the ambient $\mathscr{G}$-action of Remark 5.3. Consequently, $k$ fixes every set/function whose transitive closure is contained in $V(T)\cup\mathrm{Ord}$; in particular, if $f:X\to Y$ with $X,Y\subseteq T$, or $f:S\times\eta\twoheadrightarrow\lambda$ with $\eta,\lambda\in\mathrm{Ord}$, then $k\cdot f=f$.
@@
Put $\sigma:=\pi^{0}_{\alpha}(k)\in\mathscr{G}=\operatorname{Sym}(\omega_{1})$. Since $k\in\ker(\pi^{0}_{\alpha})$ we have $\sigma=\operatorname{id}$. By Remark 5.3, the ambient action on $A$ and on $S=A^{\omega}$ is induced from the action $\alpha\mapsto\sigma(\alpha)$ on indices of the Cohen reals, and the induced action on $\mathcal{P}(S)$ restricts to an action on the fixed subset $T=\mathcal{P}(S)^{\mathcal{N}}$. Therefore $\sigma=\operatorname{id}$ implies that $k$ fixes $A$, and hence also $S$ and $T$, pointwise.
@@
Lemma 5.6 (Orbit packages are countably closed and preserve $\mathsf{DC}$).
Work in $M\models\mathsf{ZF}+\mathsf{DC}$. Let $f\in M$ be a surjection instance and let $\mathbb{S}_{[f]}$ be its orbit package. Then:
  1. 1.
    $\mathbb{S}_{[f]}$ is countably closed;
  2. 2.
    forcing with $\mathbb{S}_{[f]}$ adds no new $\omega$-sequences of ordinals;
  3. 3.
    $M^{\mathbb{S}_{[f]}}\models\mathsf{DC}$.
@@
Let $\langle\vec{p}_{n}:n<\omega\rangle$ be a descending sequence in $\mathbb{S}_{[f]}$. Let $C:=\bigcup_{n<\omega}\operatorname{supp}(\vec{p}_{n})\subseteq[f]$. Each $\operatorname{supp}(\vec{p}_{n})$ is countable and $M\models\mathsf{DC}$, so $C$ is countable.
For each $g\in C$, consider the descending sequence $\langle\vec{p}_{n}(g):n<\omega\rangle$ in the factor $\mathbb{S}_{g}$ (interpreting $\vec{p}_{n}(g)$ as the top condition whenever $g\notin\operatorname{supp}(\vec{p}_{n})$). By Lemma 4.13 and Lemma 4.25, each factor $\mathbb{S}_{g}$ is countably closed, so choose $q_{g}\in\mathbb{S}_{g}$ with $q_{g}\leq\vec{p}_{n}(g)$ for all $n$. Define $\vec{q}$ by $\operatorname{supp}(\vec{q})=C$ and $\vec{q}(g)=q_{g}$. Then $\vec{q}\in\mathbb{S}_{[f]}$ and $\vec{q}\leq\vec{p}_{n}$ for all $n$, proving countable closure.
Items (2) and (3) follow from Lemma 4.16. ∎
@@
The countable completeness of the diagonal stage filters is what allows the CSI iteration theorem to handle limit stages. Concretely, the limit-step existence theorem for iterations of symmetric systems requires $\omega_{1}$-complete filters (CSI, [5, Theorem 3.13]), and we have imported this requirement as item (c) in Remark 2.4.
@@
Lemma 5.9 (Countable $\mathscr{G}$-support of successor coordinates).
Let $\alpha$ be a stage and let $d\in\mathrm{Coords}_{\alpha+1}\setminus\mathrm{Coords}_{\alpha}$ be a new package coordinate at the successor stage $\alpha+1$ (Definition 5.24). Then there exists a countable $E_{d}\subseteq\omega_{1}$ such that $\operatorname{Fix}(E_{d})\leq\operatorname{Stab}_{\mathscr{G}}(d)$.
@@
Write $d=\langle\alpha,i,g\rangle$ with $i<m_{\alpha}$ and $g\in[f_{\alpha,i}]$.
By the inductive hypothesis at stage $\alpha$ (see Remark 5.20), there is a countable admissible pair $(E_{f},D_{f})$ such that $\Delta^{\uparrow}_{\alpha}(E_{f},D_{f})\leq\operatorname{Stab}_{\mathscr{G}_{\alpha}}(\dot{f})$, where $\dot{f}$ is the canonical $\operatorname{HS}_{\alpha}$-name for $f$. Fix $\pi\in\operatorname{Fix}(E_{f})$. By Lemma 5.45, the diagonal lift $\widehat{\pi}^{\,D_{f}}\in\mathscr{G}_{\alpha}$ is defined, and by the choice of $(E_{f},D_{f})$ it fixes $\dot{f}$. Let $\widehat{\pi}^{\,\varnothing}\in\mathscr{G}_{\alpha}$ denote the standard lift. Then
$k:=\bigl(\widehat{\pi}^{\,\varnothing}\bigr)^{-1}\cdot\widehat{\pi}^{\,D_{f}}\in\ker(\pi^{0}_{\alpha}),$
so by Lemma 5.4 the element $k$ fixes $f$ (since $\operatorname{tc}(f)\subseteq V(T)\cup\mathrm{Ord}$). As $\widehat{\pi}^{\,D_{f}}$ fixes $f$, it follows that $\widehat{\pi}^{\,\varnothing}$ fixes $f$ as well, i.e. $\pi\cdot f=f$. Hence $\operatorname{Fix}(E_{f})\leq\operatorname{Stab}_{\mathscr{G}}(f)$.
Now if $g=\sigma\cdot f$ for some $\sigma\in\mathscr{G}$, then $\operatorname{Fix}(\sigma``E_{f})\leq\operatorname{Stab}_{\mathscr{G}}(g)$. Taking $E_{d}:=\sigma``E_{f}$ yields the desired countable support for $d$. ∎
@@
The proof uses only stage-$\alpha$ information: the inductive hypothesis that every relevant $\operatorname{HS}_{\alpha}$-name has a countable $\Delta$-support (Remark 5.20), together with the existence of diagonal lifts in $\mathscr{G}_{\alpha}$. No stage-$\alpha+1$ properties are used.
@@
Let $\mathscr{G}_{[f]}\leq\operatorname{Aut}(\mathbb{S}_{[f]})$ be the subgroup generated by all diagonal-cancellation maps $\widehat{\pi}^{\,D}$ (equivalently, $\mathscr{G}_{[f]}=\langle\Delta(E,D):(E,D)\text{ admissible}\rangle$).